OpenStudy (anonymous):

Find the zeros of the polynomial function and state the multiplicity of each. f(x) = 5(x + 8)2(x - 8)3 A. 4, multiplicity 1; -8, multiplicity 3; 8, multiplicity 3 B. -8, multiplicity 2; 8, multiplicity 3 C. -8, multiplicity 3; 8, multiplicity 2 D. 4, multiplicity 1; 8, multiplicity 1; -8, multiplicity 1

OpenStudy (whpalmer4):

Is that \[f(x) = 5(x+8)^2(x-8)^3\]?

OpenStudy (anonymous):

yes

OpenStudy (whpalmer4):

If so, the zeros of the polynomial are found by setting the products = 0. The multiplicity of each zero will just be the power of the product term in the polynomial. For example, if we had the polynomial \[f(x) = x(x-1)^3(x+2)^2\] We would have zeros at \(x = 0\), \(x=1\) and \(x=-2\), and the multiplicity would be 1, 3, and 2 respectively.

OpenStudy (anonymous):

so it will be A : 4, multiplicity 1; -8, multiplicity 3; 8, multiplicity 3